Find the slope of the following two lines:
I don't know how to do graphs so I will just explain what they look like first graph is runs horizontal on the y=-2 and the second graph runs vertical on the x=-3 the are both lines that run along that line and does not vary. are those my slopes? I am confused about this question.

Answer by stanbon(75887)

Slope of y=-2
This line passes thru all points where y=-2
Two of the points are (0,-2) and (3,-2)
The slope = [-2--2]/[3-0] = 0/3 = 0
The slope of every horizontal line is zero.
-------------------
Slope of x = -3
This line passes thru all points where x = -3
Two of the points are (-3,0) and (-3,4)
The slope = [4-0]/[-3--3] = 4/0 which is undefined.
The slope of every vertical line is undefined.
======================
Cheers,
Stan H.